find all of the possible solutions for x that satisfy the following inequality |-(x+2)^2 +1| greater than or equal to 1.

Answer by stanbon(75887)

|-(x+2)^2 +1| greater than or equal to 1
-1 <= -(x+2)^2+1 <= 1
Subtract 1 along the line to get:
-2 <= -(x+2)^2 <= 0
---
Multiply thru by -1 to get:
0 <= (x+2)^2 <= 2
---
Take the square root along the line to get:
0 <= x+2 <= sqrt(2)
---
Subtract 2 along the line to get:
-2 <= x <= sqrt(2)-2
============================
Cheers,
Stan H.